hello, i need help on solving this type of word problem, but im not sure on how to solve it the problem is Mr.Klinker is 35 and his daughter is 10. In how many years will Mr. Klinker be twice as old as his daughter? Could you please help? thank you.
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Mr.Klinker is 35 and his daughter is 10.
In how many years will Mr. Klinker be twice as old as his daughter?
-------------------
Let the number of years be "x".
Equation:
(35+x) = 2(10+x)
35 + x = 20 + 2x
x = 35-20
x = 15 (K will be twice as old as his daughter in 15 years)
====================================
Cheers,
Stan H.

Answer by Alan3354(69443)

The age difference is 25, so the daughter will be 25 and he will be 50.
Can you take it from there?